Circular No. 6018

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

LQ SAGITTARII P. Zhao and J. E. McClintock, Smithsonian Astrophysical Observa-

tory; and A. U. Landolt, Louisiana State University, write: “LQ Sgr, a poorly known nova, is now in outburst; it is ∼ 5 mag brighter in B than it was during quiescence. Its spectrum, obtained at the Cerro Tololo Interamerican Observatory (CTIO) 4-m telescope on July 2.3 UT, is that of an M giant. LQ Sgr is probably a symbiotic nova (i.e., RR-Tel class). During its eruption in 1897, it reached B = 13. Observations with the CTIO 1-m and 0.91-m telescopes give the following magnitudes and colors: July 5.23, V = 14.33, B − V = +1.57, U − B = +0.75; 6.24, V = 14.38, B − V = +1.55, U − B = +0.85, V − R = +1.98, R − I = +2.69, V − I = +4.67; 8.20, V = 14.40, B − V = +1.56, U − B = +0.87.”

SUPERNOVA 1994U IN NGC 4948 A. Boattini and M. Tombelli, Piazzano Observatory, Florence, report

the following precise position obtained on July 5.90 UT: α = 13^{h}02^{m}19^{s}.85, δ = −7^{o}40′47′′.8 (equinox 1950.0).

COMET TAKAMIZAWA (1994i)

Ephemeris from orbital elements on MPC 23652:

1994 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
July 17 11^{h}56.^{m}78 −14^{o}51.^{′}3 2.027 1.960 71.^{o}6 29.^{o}5 11.5
27 11 49.08 −15 45.5 2.252 1.977 61.3 26.8 11.7
Aug. 6 11 44.43 −16 46.1 2.460 2.001 52.0 23.5 12.0
16 11 41.84 −17 53.4 2.644 2.033 43.5 20.0 12.2
26 11 40.60 −19 07.4 2.802 2.071 35.9 16.6 12.4
Sept. 5 11 40.19 −20 28.4 2.932 2.115 29.6 13.6 12.6

Total visual magnitude estimates (cf. IAUC 6001): June 3.89 UT, 9.8 (L. Szentaskó, Veresegyház, Hungary, 0.33-m reflector); 5.31, 10.6 (A. Hale, near Orogrande, NM, 0.41-m reflector); 7.90, 9.6 (W. Hasubick, Buchloe, Germany, 25×100 binoculars); 8.87, 9.9 (G. Bakos, Ráktanya, Hungary, 0.44-m reflector); July 8.17, 10.4 (P. Roques, Williams, AZ, 0.25-m reflector + CCD).

1994 July 12 Daniel W. E. Green